OpenStudy (anonymous):

guys help me with this plz 1/a +1/b = 1/c solve for b

OpenStudy (anonymous):

what is your first step?

OpenStudy (anonymous):

i believe moving a to c

OpenStudy (anonymous):

i have the answer but i get in trouble on here for just giving answers out.

OpenStudy (anonymous):

You can get rid of the denominators by multiplying each side by a*b*c

OpenStudy (anonymous):

i got 1/b = 1/c - 1/a

OpenStudy (anonymous):

then what should i do next ?

OpenStudy (turingtest):

do what robtobey said

OpenStudy (anonymous):

If a c + b c =a b were true, could you now solve for c?

OpenStudy (anonymous):

idk :| im suck at math

OpenStudy (anonymous):

so i've got bc+ac=ab

OpenStudy (anonymous):

thats not right

OpenStudy (turingtest):

\[\frac1a+\frac1b=\frac1c\]\[\cancel abc(\frac1{\cancel a})+a\cancel bc(\frac1{\cancel b})=ab\cancel c(\frac1{\cancel c})\]\[bc+ac=ab\]

OpenStudy (turingtest):

now get all the terms that have b on the same side of the equals sign, then factor out the b

OpenStudy (anonymous):

so the final answer is \[\frac{ ac }{a-c }\]

OpenStudy (turingtest):

yeah looks good :)

OpenStudy (anonymous):

tks :)